HRMS, 28th Colloquium, Dijon, France, Aug. 28-Sep. 1, 2023

ANHARMONIC AND CORIOLIS INTERACTIONS IN THE v_{3} = 2/v_{2} = 1/v_{5} = 1/v_{3} = v_{6} = 1 LEVEL SYSTEM OF CH_{3}Br

\underline{ADINA CEAUSU-VELCESCU} , Université de Perpignan, LAMPS, 52 Avenue Paul Alduy, 66860 Perpignan Cedex, France; WILFRIED TCHANA BETNGA,

\underline{FRIDOLIN KWABIA TCHANA} , Université Paris-Cité and Univ. Paris Est Créteil, CNRS, LISA, F-75013 Paris, France; \underline{LAURENT MANCERON} , Ligne

AILES, Synchrotron Soleil, L’Orme des Merisiers, St-Aubin, BP48, 91192 Gif-sur-Yvette, France; Université Paris-Cité and Univ. Paris Est Créteil, CNRS,

LISA, F-75013 Paris, France

The ν_{2}/ν_{5} band system of CH_{3}Br, a methyl halide playing an important role in the catalytic destruction of stratospheric ozone, already constituted the subject of several medium and high-resolution studies. The novelty of the present study consists in considering, besides the strong Coriolis and α-interactions coupling the v_{2} = 1 and v_{5} = 1 levels, a large variety of anharmonic and rovibrational interactions involving also the v_{3} = 2 and v_{3} = v_{6} = 1 levels. Thousands of new data, belonging either to high J and K values in the ν_{2} and ν_{5} bands or to the, very weak, ν_{3} + ν_{6} combination band, were included in the least-squares calculations.

Thanks to the large set of data, including more than 6300 experimental wavenumbers of the ν_{2}, ν_{5}, 2ν_{3} and ν_{3} + ν_{6} rovibrational bands, with J ≤ 74 and K ≤ 18, combined to the completeness of the theoretical model, the global standard deviation,

p-number: p062  Submitted on Fri, 09 Jun 2023 17:16:50 +0200

of 2.34 × 10^{−4} cm^{−1}, represents a great improvement with respect to the previous high-resolution study of the ν_{2}/ν_{5} band system^{1}. Moreover, the present study explores also the reductions’ issue, in the spirit of the work of Střiteská et al^{2} and Sarka et al^{3}. Two different reductions schemes were thus applied and were proved to be equally successful.

CH_{3}^{79}Br:

Parameter  QQ  QC
(η_{J}^{5} + η_{K}^{5}) × 10^{5} -20.9766(47) -20.9704(47)
α_{2}^{A} × 10^{3}  -21.8459(36) -21.8479(36)
α_{5}^{A} × 10^{3}  46.7461(36) 46.7461(36)
(α_{2}^{B} + 4q_{22}) × 10^{3} 1.02914(10) 1.02865(11)
(α_{5}^{B} − 2q_{22}) × 10^{3} -0.032513(23) -0.031836(69)
(α_{2}^{B} + 2α_{5}^{B}) × 10^{3} 0.96412(11) 0.964981(95)

CH_{3}^{81}Br:

Parameter  QQ  QC
(η_{J}^{5} + η_{K}^{5}) × 10^{5} -21.0762(71) -21.0639(71)
α_{2}^{A} × 10^{3}  -21.8565(43) -21.8587(44)
α_{5}^{A} × 10^{3}  46.7752(43) 46.7840(43)
(α_{2}^{B} + 4q_{22}) × 10^{3} 1.02346(22) 1.02285(22)
(α_{5}^{B} − 2q_{22}) × 10^{3} -0.03028(13) -0.02952(13)
(α_{2}^{B} + 2α_{5}^{B}) × 10^{3} 0.96290(34) 0.96382(14)

1F. Kwabia Tchana, I. Kleiner, J. Orphal, N. Lacome, O. Bouba, J. Mol. Spectrosc. 228 (2004)

441-452.

2L. Nová Střiteská, K. Sarka, Š Urban, J. Mol. Spectrosc. 256 (2009) 135-140.

3K. Sarka, L. Nová Střiteská, A. Ceausu-Velcescu, J. Mol. Spectrosc. 311 (2015) 84-99.

p-number: p062  Submitted on Fri, 09 Jun 2023 17:16:50 +0200